# Kernel-Based Nonparametric Tests For Shape Constraints

### v1

@@ 1.2. Contributions
    We propose a Wald-type test statistic for shape restrictions over a finite grid (positivity, monotonicity, convexity, etc.). The test statistic is a Mahalanobis distance and admits an implementation based on a non-negative least squares program.
@@
    We provide an efficient computation procedure that can be applicable for large datasets, with large samples, and can be used for testing on dense grids.
@@ 2.1. Notation and setting
We define the set $\mathcal{A}_{s}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathbb{N}^{d}:|\boldsymbol{\alpha}|\leq s\}$ and $m_{s}\mathrel{\mathrel{\mathop{:}}=}|\mathcal{A}_{s}|=\binom{s+d}{d}$. In addition, we utilize the usual notions of $o_{\mathbb{P}}$ and $\mathcal{O}_{\mathbb{P}}$, and refer the reader to Vaart (1998) for details.
@@ 2.3.1. Population and empirical problem
In many decision problems, the target object is a linear score that depends on the level and on the gradient information of an unknown function $h$. A generic specification is given by (2.3), where $\boldsymbol{z}=(\boldsymbol{x},y)$ denotes observed data and the weights $w_{\boldsymbol{\alpha}}(\cdot)$ encode the task via the dependencies on $h$ and its gradients. In the case, the target functional only depends on the gradient values/specified derivatives only, we can consider $\mathcal{A}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathcal{A}_{s}:w_{\boldsymbol{\alpha}}\equiv 0\}$. Such a form allows for flexibility and covers, for example, portfolio rules or control scores that use the function value as a signal and gradient or curvature components as sensitivity adjustments. Since each $D^{\boldsymbol{\alpha}}h(\boldsymbol{x})$ is a bounded linear functional in an RKHS with a sufficiently smooth kernel, $\mathcal{R}(h;\boldsymbol{z})$ remains linear in $h$ and admits a representer $\psi(\boldsymbol{z})$.
@@ 4.1.3. Test statistic
Theorem 4.2 shows the consistency of the finite-sample $n\times n$ covariance matrix $\widehat{\boldsymbol{\Omega}}_{\lambda}$. As a final step towards constructing the test statistic, we define the following. We refer an interested reader to Silvapulle and Sen (2001, Chapter 3.4) for further details.
@@
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a closed, convex cone, and let $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{V})$ where $\boldsymbol{V}$ is a positive definite matrix. Then, $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is defined to be the random variable having the same distribution as
@@
Under the inner product $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\boldsymbol{V}^{-1}}=\boldsymbol{x}^{\top}\boldsymbol{V}^{-1}\boldsymbol{y}$, (4.11) may be written as, see Silvapulle and Sen (2001, Proposition 3.4.1)
$\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})\sim\|\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})\|_{\boldsymbol{V}^{-1}}^{2},$
where $\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})$ is the orthogonal projection of $\boldsymbol{Z}$ onto $\mathcal{M}$ under the inner product $\langle\cdot,\cdot\rangle_{\boldsymbol{V}^{-1}}$. We also have the following result.
@@
Let $\mathcal{M}$ be a closed, convex cone in $\mathbb{R}^{n}$ and let $\boldsymbol{V}\in\mathbb{R}^{n\times n}$ be a symmetric and positive definite matrix. Then the distribution of $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is given by
(4.12) $\mathbb{P}\left(\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})\leq c\right)=\sum_{j=0}^{n}w_{j}\,\mathbb{P}(\chi^{2}_{j}\leq c),$
where $w_{j}\geq 0$ for $0\leq j\leq n$ and $\sum_{j=0}^{n}w_{j}=1$.
Remark 4.5.
From the projection theorem of Bauschke and Combettes (2017, Theorem 3.16), if $\mathcal{M}$ is a non-empty closed convex subset of a Hilbert space $\mathscr{H}$, then for any $u\in\mathscr{H}$, the orthogonal projection $\Pi_{\mathcal{M}}(u)$ is well-defined and unique and satisfies
@@
We now state the following theorem, which defines the test statistic and shows its asymptotic distribution, see Appendix C for a proof.
@@
(4.15) $W_{N}\overset{d}{\longrightarrow}W\sim\chi_{n}^{2}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+}).$
For a fixed derivative order $\boldsymbol{\alpha}\in\mathcal{A}_{s}$ and a grid $\mathcal{G}\subset\mathcal{X}$, we test the positivity constraint of the $\boldsymbol{\alpha}$ derivative evaluation at the grid points by setting up the hypothesis test as:
(4.16) $H_{0}:\boldsymbol{\theta}\in\mathbb{R}^{n}_{+}\quad\text{vs.}\quad H_{1}:\,\text{there exists some $j\in\{1,\ldots,n\}$ such that $h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})<0$}.$
The least favorable null is given by $\boldsymbol{\theta}=\boldsymbol{0}$, that is, the vector of the derivative evaluations at the grid points given by $\boldsymbol{\theta}$ lies on the boundary of the positive orthant $\mathbb{R}^{n}_{+}$. Under this, the Wald-type statistic given by $W_{N}$ is the distance of the centered, scaled estimate $\widehat{\boldsymbol{\theta}}$ to the closed, convex cone given by the positive orthant under the Mahalanobis distance defined in Definition 4.3. This can be solved as a non-negative least squares problem, see more details in Appendix E.
@@ Appendix B Lemmas for Section 4
Thus, $\boldsymbol{u}_{k}$ is bounded in the Euclidean norm and hence, by the Bolzano-Weierstrass theorem, there exists a convergent subsequence $\boldsymbol{u}_{k_{\ell}}\to{\boldsymbol{u}}^{\ast}$, where we consider the convergence in the topology induced by the usual Euclidean norm. From Remark 4.5, the projection $\boldsymbol{u}_{k_{\ell}}$ uniquely satisfies
@@
But this is the inequality characterizing the unique projection $\boldsymbol{u}^{\ast}=\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})$, see Remark 4.5. Hence, the set of subsequential limits of $\boldsymbol{u}_{k}$ is unique and any convergent subsequence of $\{\boldsymbol{u}_{k}\}_{k\in\mathbb{N}}$ has the same limit $\boldsymbol{u}^{\ast}$. This property and the fact that $\{\boldsymbol{u}_{k}\}$ is bounded in the Euclidean norm imply that the sequence $\boldsymbol{u}_{k}$ converges to the limit $\boldsymbol{u}^{\ast}$. Since $\boldsymbol{x}\in\mathbb{R}^{n}$ is arbitrary, therefore,
@@
Hence, we have from Vaart [1998, Theorem 2.7], for asymptotically large $N$,
$\left(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\right)\overset{d}{\longrightarrow}\left(\boldsymbol{Z},\boldsymbol{\Omega}_{\lambda}^{-1}\right)\quad\text{on}\quad\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}.$
Since $\mathbb{R}^{n}_{+}$ is a closed convex cone, $N\mathbb{R}^{n}_{+}=\mathbb{R}^{n}_{+}$ for any $N\geq 0$. Hence,
@@
From Lemma C.3, $W_{N}$ is continuous as a function of $\left(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\right)$. Hence, by the CMT,
@@
Since $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{\Omega}_{\lambda})$, hence $\boldsymbol{Z}^{\top}\boldsymbol{\Omega}_{\lambda}^{-1}\boldsymbol{Z}\sim\chi^{2}_{n}$. Hence, from Definition 4.3, it follows that
$W=\|\boldsymbol{Z}-\Pi^{\boldsymbol{\Omega}_{\lambda}^{-1}}_{\mathbb{R}^{n}_{+}}(\boldsymbol{Z})\|^{2}_{{\boldsymbol{\Omega}}_{\lambda}^{-1}}\sim\chi^{2}_{n}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+}).$

### v2

@@ 1.2. Contributions
    We propose a Wald-type test statistic for assessing shape restrictions over a finite grid (positivity, monotonicity, convexity, etc.). The test statistic measures a squared Mahalanobis distance of the projection error admits an implementation based on a non-negative least squares program.
@@
    We provide an efficient computation procedure based on a pivoted Cholesky decomposition that can be handle large datasets, with large samples, and can be used for testing on dense grids.
@@ 2.1. Notation and setting
We define the set $\mathcal{A}_{s}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathbb{N}^{d}:|\boldsymbol{\alpha}|\leq s\}$ and $m_{s}\mathrel{\mathrel{\mathop{:}}=}|\mathcal{A}_{s}|=\binom{s+d}{d}$. In addition, we utilize the usual notions of $o_{\mathbb{P}}$ and $\mathcal{O}_{\mathbb{P}}$, and refer the reader to van der Vaart (1998) for details.
@@ 2.3.1. Population and empirical problem
In many decision problems, the target object is a score that depends linearly on the level and on the gradient information of an unknown function $h$. A generic specification is given by (2.3), where $\boldsymbol{z}=(\boldsymbol{x},y)$ denotes observed data and the weights $w_{\boldsymbol{\alpha}}(\cdot)$ encode the task via the dependencies on $h$ and its gradients. In the case, the target functional only depends on the gradient values/specified derivatives only, we can consider $\mathcal{A}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathcal{A}_{s}:w_{\boldsymbol{\alpha}}\equiv 0\}$. Such a form allows for flexibility and covers, for example, portfolio rules or control scores that use the function value as a signal and gradient or curvature components as sensitivity adjustments. Since each $D^{\boldsymbol{\alpha}}h(\boldsymbol{x})$ is a bounded linear functional in an RKHS with a sufficiently smooth kernel, $\mathcal{R}(h;\boldsymbol{z})$ remains linear in $h$ and admits a representer $\psi(\boldsymbol{z})$.
@@ 4.1.3. Test statistic
Theorem 4.2 shows the consistency of the finite-sample $n\times n$ covariance matrix $\widehat{\boldsymbol{\Omega}}_{\lambda}$. As a final step towards constructing the test statistic, we define the following. We refer an interested reader to Silvapulle and Sen (2001) for further details.
@@
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a closed convex cone, and let $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{V})$ where $\boldsymbol{V}$ is a symmetric and positive definite matrix. Then, $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is defined to be the random variable having the same distribution as
@@
Denote by $\mathcal{M}^{\circ}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{x}:\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\boldsymbol{V}^{-1}}\leq 0\text{ for all $\boldsymbol{y}\in\mathcal{M}$}\}$ the polar cone of $\mathcal{M}$, where we define the inner product $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\boldsymbol{V}^{-1}}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{x}^{\top}\boldsymbol{V}^{-1}\boldsymbol{y}$. From Silvapulle and Sen (2001, Proposition 3.4.1), we have:
(4.12) $\|\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})\|_{\boldsymbol{V}^{-1}}^{2}\sim\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M}),\qquad\|\boldsymbol{Z}-\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})\|_{\boldsymbol{V}^{-1}}^{2}\sim\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M}^{\circ}),$
where $\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})$ is the orthogonal projection of $\boldsymbol{Z}$ onto $\mathcal{M}$ under the inner product $\langle\cdot,\cdot\rangle_{\boldsymbol{V}^{-1}}$. Moreover, we have:
@@
Let $\mathcal{M}$ be a closed convex cone in $\mathbb{R}^{n}$ and let $\boldsymbol{V}\in\mathbb{R}^{n\times n}$ be a symmetric and positive definite matrix. Then the distribution of $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is given by
(4.13) $\mathbb{P}\left(\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})\leq c\right)=\sum_{j=0}^{n}w_{j}(n,\boldsymbol{V},\mathcal{M})\,\mathbb{P}(\chi^{2}_{j}\leq c),$
where $w_{j}(n,\boldsymbol{V},\mathcal{M})\geq 0$ for $0\leq j\leq n$ and $\sum_{j=0}^{n}w_{j}(n,\boldsymbol{V},\mathcal{M})=1$.
We also put the following result, which characterizes an orthogonal projection in Hilbert spaces.
Theorem 4.5 (Bauschke and Combettes (2017, Theorem 3.16)).
Let $\mathcal{M}$ be a non-empty closed convex subset of a Hilbert space $\mathscr{H}$. Then for any $u\in\mathscr{H}$, the orthogonal projection $\Pi_{\mathcal{M}}(u)$ (under the $\mathscr{H}$-inner product) is well-defined and unique and satisfies
@@
For a fixed derivative order $\boldsymbol{\alpha}\in\mathcal{A}_{s}$ and a grid $\mathcal{G}\subset\mathcal{X}$, we test the one-sided composite cone restriction given by the positivity constraint of the $\boldsymbol{\alpha}$-derivative evaluation at the grid points:
(4.15) $H_{0}:\boldsymbol{\theta}=[h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})]_{j=1}^{n}\in\mathbb{R}^{n}_{+}\;\text{vs.}\;H_{1}:\,\text{there exists some $j\in\{1,\ldots,n\}$ such that $h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})<0$}.$
The least favorable null is given by $\boldsymbol{\theta}=\boldsymbol{0}$, that is, the boundary of the positive orthant $\mathbb{R}^{n}_{+}$. We now state the following theorem, which defines the test statistic and shows its asymptotic distribution, see Appendix C for a proof.
@@
(4.17) $W_{N}\overset{d}{\longrightarrow}W\sim\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\bigl(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi_{n}^{2}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+}).$
Moreover, we have $\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi^{2}_{n}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})$, where the equality holds almost surely.
Under the least favorable null, the Wald-type statistic given by $W_{N}$ is the distance of the centered, scaled estimate $\sqrt{N}\widehat{\boldsymbol{\theta}}$ to the closed, convex cone given by the positive orthant $\mathbb{R}^{n}_{+}$. Tests for the opposite sign (e.g., monotonically decreasing or concavity) are obtained by applying the positivity test to the sign-flipped vector $\boldsymbol{\theta}$.
Remark 4.7.
Unlike the usual form of the Wald test, here we have a one-sided test. The test statistic $W_{N}$ measures the projection error under the Mahalanobis distance $\|\cdot\|_{\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}}$, and its limit law describes how far $\boldsymbol{\theta}$ is from the feasibility region, which is given by the composite null hypothesis. Under $H_{0}$, the asymptotic distribution of $W_{N}$ depends on which inequalities are binding for $\boldsymbol{\theta}$. As the entries get more strictly positive (when $\boldsymbol{\theta}$ moves into the interior of $\mathbb{R}^{n}_{+}$), the test statistic $W_{N}$ gets stochastically smaller: the largest (least favorable case) occurs when all the constraints are binding, that is, all the entries of $\boldsymbol{\theta}$ are zero. In particular, for any $\boldsymbol{\theta}\in\mathbb{R}^{n}_{+}$, $\mathbb{P}_{\boldsymbol{\theta}}(W_{N}\geq c)\leq\mathbb{P}_{\boldsymbol{\theta}=\boldsymbol{0}}(W_{N}\geq c)$. So, we calibrate the critical values (or $p$-values) at the least favorable null $\boldsymbol{\theta}=\boldsymbol{0}$.
The asymptotic distribution of $W_{N}$ stated in Theorem 4.6 has the form as given in Theorem 4.4 and hence, to obtain the $p$-values, we need to calculate $\mathbb{P}\left(\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})\leq c\right)$. In practice, the tail-probability is estimated via a Monte Carlo replication, see Silvapulle and Sen (2001, Section 3.5), such that the test statistic is solved via a non-negative least squares problem, see Appendix E.
@@ Appendix B Lemmas for Section 4
Thus, $\boldsymbol{u}_{k}$ is bounded in the Euclidean norm and hence, by the Bolzano-Weierstrass theorem, there exists a convergent subsequence $\boldsymbol{u}_{k_{\ell}}\to{\boldsymbol{u}}^{\ast}$, where we consider the convergence in the topology induced by the usual Euclidean norm. From Theorem 4.5, the projection $\boldsymbol{u}_{k_{\ell}}$ uniquely satisfies
@@
But this is the inequality characterizing the unique projection $\boldsymbol{u}^{\ast}=\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})$, see Theorem 4.5. Hence, the set of subsequential limits of $\boldsymbol{u}_{k}$ is unique and any convergent subsequence of $\{\boldsymbol{u}_{k}\}_{k\in\mathbb{N}}$ has the same limit $\boldsymbol{u}^{\ast}$. This property and the fact that $\{\boldsymbol{u}_{k}\}$ is bounded in the Euclidean norm imply that the sequence $\boldsymbol{u}_{k}$ converges to the limit $\boldsymbol{u}^{\ast}$. Since $\boldsymbol{x}\in\mathbb{R}^{n}$ is arbitrary, therefore,
@@
Hence, we have from van der Vaart [1998, Theorem 2.7], for asymptotically large $N$,
$\bigl(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\bigr)\overset{d}{\longrightarrow}\bigl(\boldsymbol{Z},\boldsymbol{\Omega}_{\lambda}^{-1}\bigr)\quad\text{on}\quad\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}.$
Since $\mathbb{R}^{n}_{+}$ is a closed convex cone, $\sqrt{N}\mathbb{R}^{n}_{+}=\mathbb{R}^{n}_{+}$ for any $N\geq 1$. Hence,
@@
From Lemma C.3, $W_{N}$ is continuous as a function of $\bigl(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\bigr)$. Hence, by the CMT,
@@
From (4.12), it follows, $W\sim\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})$. Moreover, from Moreau’s decomposition and the Pythagorean identity,
$\boldsymbol{Z}^{\top}{\boldsymbol{\Omega}}_{\lambda}^{-1}\boldsymbol{Z}=\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})+\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})$
Since $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{\Omega}_{\lambda})$, hence $\boldsymbol{Z}^{\top}\boldsymbol{\Omega}_{\lambda}^{-1}\boldsymbol{Z}\sim\chi^{2}_{n}$. Thus, $\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi^{2}_{n}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})$, where the equality holds almost surely. ∎